Calcultator Units-of-activity Depreciation A truck acquired at a cost of $375,000 has an estimated residual value of $21,600, has an estimated useful life of 38,000 miles, and was driven 2,700 miles during the year. Determine the following. If required, round your answer for the depreclation rate to two decimal places (a) The depreciable cost (b) The depreciation rate per mile degpreciation for the year (c) The units-of-activity depreciation for the year
invoker- Show Me HowCalculator Stralight-Line Depreciation A building acquired at the beginning of the year at a cost of $80,100 has an estimated residual value of $3,100 and an estimated useful life of 10 years. Determine the following: (a) The depreciable cost (b) The straight-Ine rate (c) The annual straight-line depreciation
Answer #1

(In $)

A) (a) Depreciable cost = Cost - salvage value

= 375,000-21,600

= 353,400

(b) Depreciation rate per mile = Depreciable cost/ useful life

= 353,400/38,000

= 9.30 per mile

(c) units of activity depreciation per the year = annual units × depreciation per unit

= 2,700× 9.30

= 25,110

B) (a) Depreciable cost = Cost - Residual value

= 80,100-3,100

= 77,000

(b) Straight line rate = Straight line Depreciation/ Depreciable cost

= 7,700/77,000

= 10%

(c) Annual straight line depreciation = Depreciable cost / Estimated useful life

= 77,000/10

= 7,700